### v1

@@ 3.1 Nonlocal Poisson equation
Our approach for solving (8) builds on the spectral method for nonlocal Poisson equation on periodic domains by constructing a periodic extension of the problem using the Fourier continuation algorithm. We then solve the extended problem using the iterative linear algebra solver GMRES [32] to obtain the solution within a specified tolerance. This process is detailed below.
@@
$Au=g,\quad A=\begin{pmatrix}L\\ B\end{pmatrix},\quad g=\begin{pmatrix}f\\ b\end{pmatrix},$ (9)
where $L$ and $B$ are respectively the associated matrices of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ and the boundary operator $\mathcal{B}$ given by the Dirichlet conditions on the collar region. To apply GMRES to this system, we do not need to compute $A$ and store it in memory. Instead, it is sufficient to express the forward operator $v\mapsto Av$ as a function. The action of the operator $\mathcal{B}$ is a restriction of the values of $v$ to the collar region. We derive the forward operator function using the properties of $\mathcal{L}^{\delta,\beta}$ as follows:
@@
where $u_{\text{ext}}$ is the extension of $u$ obtained by the 2D-FC method, $\widehat{u}_{\text{ext}}$ is the Fourier transform of $u_{\text{ext}}$ obtained by means of the fast Fourier transform (FFT), $Lu_{\text{ext}}$ is the inverse Fourier transform of $\widehat{Lu_{\text{ext}}}$, and $Lu$ is the restriction of $Lu_{\text{ext}}$ to the domain $\Omega$. We then obtain the solution of (8) by applying GMRES to the system (9) within a specified tolerance.
@@ 3.2 Nonlocal diffusion equation
$\vec{u}_{k}\to(\vec{u}_{k})_{\text{ext}}\to\widehat{(\vec{u}_{k})}_{\text{ext}}\to\widehat{(\mathcal{L}^{\delta,\beta}\vec{u}_{k})}_{\text{ext}}=m^{\delta,\beta}\widehat{(\vec{u}_{k})}_{\text{ext}}\to(\mathcal{L}^{\delta,\beta}\vec{u}_{k})_{\text{ext}}\to\mathcal{L}^{\delta,\beta}\vec{u}_{k}+\vec{s}_{k}=f(\vec{u}_{k}),$
@@ 4.1 Accuracy and convergence test for nonlocal Poisson equation
$0.4$ $-82.87098585883194$ $-180.5053934013443$ $-387.0397711705603$
Table 1: The multipliers of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ applied on $u=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $r_{1}=10.6418$ and $r_{2}=12.6418$, with a fixed $\delta=0.4$ and three different values of $\beta$.
We compute these multipliers using the implementation of the ${}_{2}F_{3}$ hypergeometric function provided by mpmath [24], as detailed in [2]. We then use these manufactured solutions to compute the errors as shown in the following example.
@@
[…]
@@
A convergence study for this test case is shown in Figure 8 for $d=4$ and $d=5$, with fixed $M=d+1$. We use the manufactured solution described above to compute the $L^{2}$ errors reported. We observe seventh and eighth orders of convergence for $d=4$ and $d=5$, respectively, showing convergence of order $d+3$ for the three cases of $\beta$ considered. Moreover, we observe no significant difference in the errors reported for $d=4$ and $d=5$ for these numbers of spatial discretizations. Thus, for the rest of this work, we will use $d=4$ for all numerical computations of nonlocal Poisson equations.
(a) $d=4$
(b) $d=5$
Figure 8: Convergence of the 2D-FC based solution of the nonlocal Poisson equation (8) in the settings of Example 3 with fixed $\delta=0.4$, fixed $M=d+1$, and two different values of $d$.
@@ 4.2 Accuracy and convergence test for nonlocal diffusion equation
A direct computation of the multipliers $m^{\delta,\beta}$ is provided in Table 2. We compute these multipliers using the implementation of the ${}_{2}F_{3}$ hypergeometric function provided by mpmath [24], as detailed in [2].
@@
$0.3$ $-130.16228859689554$ $-321.3202730766787$ $-689.8419741563309$
Table 2: The multipliers of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ applied on $u=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $r_{1}=r_{2}=15.6455$, with a fixed $\delta=0.3$ and three different values of $\beta$.
@@
[…]
using step sizes $h=0.02/2^{k},~{}k=0,1,2,3,4$ and time step size $\tau=5.0\cdot 10^{-07}$. We fix $\delta=0.3$, and choose three values of $\beta:$ $1.0$, $2.0$, and $2.5$, and we set the boundary function $b(x,y)=e^{-2\pi^{2}\kappa t}\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $\kappa=0.1$, $r_{1}=r_{2}=15.6455$, and the heat source, $s(x,y,t)$, is as given in (13).
@@
$5.00\cdot 10^{-3}$ $5.53\cdot 10^{-09}$ $7.3$ $3.30\cdot 10^{-08}$ $7.5$ $3.23\cdot 10^{-07}$ $7.0$
@@
Table 3: Convergence of the 2D-FC based solution of the nonlocal diffusion equation (11) in the settings of Example 4 with fixed $\delta=0.3$, fixed $M=d+1$, two different values of $d$: $4$ and $5$, and three choices of $\beta$: $1.0$, $2.0$, and $2.5$.
@@ 4.3 Study of discontinuities for the nonlocal Poisson equation
Let $\Omega$ be the kite-shaped domain contained within the curve defined in (14). We solve the nonlocal Poisson equation (8) over $\Omega$ using $h=0.0025$ with $\delta=0.4$, and three different values of $\beta:$ $1.0$, $2.0$, and $3.0$, and we set $f(x,y)=4$ and $b(x,y)=x^{2}+y^{2}$.
In this example, since for any $\delta>0$ and $\beta<4$, $\mathcal{L}^{\delta,\beta}(x^{2}+y^{2})=4$ for all $x\in\Omega$, then the unique solution for this problem is given explicitly by $u(x,y)=x^{2}+y^{2}$, which is obviously a continuous function. The numerical solution obtained using our solver is consistent with the above observation as illustrated in Figure 9. To emphasize, the numerical solution represent
@@
Let $\Omega$ be the unit disk. We solve the nonlocal Poisson equation (8) over $\Omega$ with $h=0.0025$. We choose $\delta=0.2$, four values of $\beta:$ $1.5$, $2.0$, $2.2$, $3.1$, and we set $f(x,y)=|\sin(\pi x)\sin(\pi y)|$ and $b(x,y)=-x^{2}$.
@@
magnitude of the $\text{magnitude of the jump}=\left|\widetilde{u}(x_{\theta},y_{\theta})-b(x_{\theta},y_{\theta})\right|,$
@@
Let $\Omega$ be the kite-shaped domain contained within the curve defined in (14). We solve the nonlocal Poisson equation (8) over $\Omega$ with $h=0.0025$. We choose $\delta=0.5$, four different values of $\beta:$ $1.2$, $2.0$, $2.5$, $3.1$, and we set the boundary function $b(x,y)=x^{2}+y^{2}$ and
@@
magnitude of the $\text{magnitude of the jump}=\left|\widetilde{u}_{\text{int}}(x_{\theta},y_{\theta})-\widetilde{u}_{\text{ext}}(x_{\theta},y_{\theta})\right|,$
@@
To make a connection between the numerical results in this work, in particular Example 7, and previous mathematical analysis in this area, we note that regularity results for the nonlocal Poisson equation on periodic domains have been developed in [2]. However, regularity of solutions for the nonlocal equation on bounded domains is still an open problem.
For the periodic case, Theorem 4.1 in [2] implies that if $f\in L^{2}$, then $u\in H^{\beta-n}$. In particular, in the case when when $n=2$ and $43<\beta<$ 4, then $u\in H^{\beta-2}$, which implies that $u$ is continuous using the embedding theorem. Using intuition from the periodic case combined with the numerical observations in Example 7, we expect that this smoothing behavior as $\beta$ increases would be true in the case of bounded domains as stated in the following conjecture.
@@
Let $u$ be the solution of (8) and assume that $f\in L^{2}(\Omega)$. Then, if $43<\beta<$ 4, then $u\in C(\Omega)$.
@@ 4.4 Evolution of discontinuities for nonlocal diffusion equation
where $r(\theta)=1.1+\cos(7\theta)/20+\sin(4\theta)/30,$ and $0\leq\theta\leq 2\pi$. We solve the nonlocal diffusion equation (11) over $\Omega$ at four different time instances: $0.02$, $0.03$, $0.04$, and $0.05$ with time step size $\tau=2.5\cdot 10^{-7}$ and spatial step size $h=0.005$. We set the initial condition $u_{0}(x,y)=\sin(2\pi x)\cos(2\pi y)$, the boundary function $b(x,y)=-(1+x^{2}+y^{2})$, and the heat source $s(x,y,t)=0.$
@@
We solve the nonlocal diffusion equation (8) over the domain contained within the curve given in (15) with spatial step size $h=0.005$ and time step size $\tau=2.5\cdot 10^{-7}$ at four different time instances: $t=0.001$, $0.005$, $0.01$, and $0.03$. We set $b(x,y)=0$, the heat source $s(x,y,t)=0$, and the initial condition
$u_{0}(x,y)=\begin{cases}2,&\text{ if }x^{2}+y^{2}<0.3,\\ e^{-x^{4}-y^{4}},&\text{ otherwise}.\end{cases}$

### v2

@@ 3.1 Nonlocal Poisson equation
Our approach for solving (9) builds on the spectral method for nonlocal Poisson equation on periodic domains by constructing a periodic extension of the problem using the Fourier continuation algorithm. We then solve the extended problem using the iterative linear algebra solver GMRES [saad2003iterative] to obtain the solution within a specified tolerance. This process is detailed below.
@@
$Au=g,\quad A=\begin{pmatrix}L\\ S\end{pmatrix},\quad g=\begin{pmatrix}f\\ b\end{pmatrix}$, (10)
where $L$ and $S$ are respectively the associated matrices of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ and the boundary operator $\mathcal{S}$ given by the Dirichlet conditions on the collar region. To apply GMRES to this system, we do not need to compute $A$ and store it in memory. Instead, it is sufficient to express the forward operator $v\mapsto Av$ as a function. The action of the operator $\mathcal{S}$ is a restriction of the values of $v$ to the collar region. We derive the forward operator function using the properties of $\mathcal{L}^{\delta,\beta}$ as follows:
@@
where $u_{\text{ext}}$ is the extension of $u$ obtained by the 2D-FC method, $\widehat{u}_{\text{ext}}$ is the Fourier transform of $u_{\text{ext}}$ obtained by means of the fast Fourier transform (FFT), $Lu_{\text{ext}}$ is the inverse Fourier transform of $\widehat{Lu_{\text{ext}}}$, and $Lu$ is the restriction of $Lu_{\text{ext}}$ to the domain $\Omega$. The extension step of (11) has a computational complexity of $\mathcal{O}(N^{\frac{1}{2}})$; each interpolation and each blending-to-zero has constant complexity. Applying the operator in Fourier space is a $\mathcal{O}(N)$ operation, and the final restriction operator is $\mathcal{O}(1)$. Thus, the total computational complexity of (11) is dominated by the $\mathcal{O}(N\log N)$ FFT. We obtain the solution of (9) by applying GMRES to the system (10) within a specified tolerance.
@@ 3.2 Nonlocal diffusion equation
$\vec{u}_{k}\to(\vec{u}_{k})_{\text{ext}}\to\widehat{(\vec{u}_{k})}_{\text{ext}}\to\widehat{(\mathcal{L}^{\delta,\beta}\vec{u}_{k})}_{\text{ext}}=m^{\delta,\beta}\widehat{(\vec{u}_{k})}_{\text{ext}}\to\mathcal{L}^{\delta,\beta}\vec{u}_{k}+\vec{s}_{k}=f(\vec{u}_{k}),$
@@ 4.1 Accuracy and convergence test for nonlocal Poisson equation
$0.4$ 82.87098585883194-82.87098585883194 180.5053934013443-180.5053934013443 387.0397711705603-387.0397711705603
Table 1: The multipliers of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ applied on the function $u=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $r_{1}=10.6418$ and $r_{2}=12.6418$, with a fixed $\delta=0.4$ and three different values of $\beta$.
We compute these multipliers using the implementation of the $_{2}F_{3}$ hypergeometric function provided by mpmath [johansson2013mpmath], as detailed in [alali2021fourier]. We then use these manufactured solutions to compute the errors as shown in the following example.
@@
We employ our proposed spectral method to solve the nonlocal Poisson equation (9) over the kite-shaped domain contained within the curve given by
and study the error convergence using step sizes $h=0.02/2^{k},~k=0,1,2,3,4$. We fix $\delta=0.4$, and choose three different values of $\beta:$ $1.2$, $2.0$, $2.5$. We set the boundary function $b(x,y)=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$ and the right-hand side function
@@
A convergence study for this test case is shown in Table 2 for $d=4$ and $d=5$, with fixed $M=d+1$. We use the manufactured solution described above to compute the $L^{2}$ errors reported. We observe sixth and seventh orders of convergence for $d=4$ and $d=5$, respectively, showing convergence of order $d+2$ for the three cases of $\beta$ considered. Moreover, we observe no significant difference in the errors reported for $d=4$ and $d=5$ for these step sizes. Thus, for the rest of this work, we will use $d=4$ for all numerical computations of nonlocal Poisson equations.
  $d$ $h$ $\overset{\beta=1.2}{\varepsilon_{2}}$ Order $\overset{\beta=2.0}{\varepsilon_{2}}$ Order $\overset{\beta=2.5}{\varepsilon_{2}}$ Order
4 $2.00\cdot 10^{-2}$ $1.29\cdot 10^{-04}$ —– $3.56\cdot 10^{-04}$ —– $8.97\cdot 10^{-04}$ —–
$1.00\cdot 10^{-2}$ $2.00\cdot 10^{-06}$ $6.0$ $4.10\cdot 10^{-06}$ $6.4$ $1.09\cdot 10^{-05}$ $6.4$
$5.00\cdot 10^{-3}$ $5\cdot 10^{-08} 6.26.2$ $7.75\cdot 10^{-08}$ $5.7$ $2.32\cdot 10^{-07}$ $5.6$
$2.50\cdot 10^{-3}$ $1.44\cdot 10^{-10}$ $7.5$ $2.19\cdot 10^{-10}$ $8.5$ $1.61\cdot 10^{-09}$ $7.2$
$1.25\cdot 10^{-3}$ $1.62\cdot 10^{-12}$ $6.5$ $4.79\cdot 10^{-12}$ $5.5$ $2.24\cdot 10^{-11}$ $6.2$
5 $2.00\cdot 10^{-2}$ $6.65\cdot 10^{-04}$ —– $2.09\cdot 10^{-03}$ —– $5.11\cdot 10^{-03}$ —–
$1.00\cdot 10^{-2}$ $6\cdot 10^{-06} 6.66.6$ $1.71\cdot 10^{-05}$ $6.9$ $4.99\cdot 10^{-05}$ $6.7$
$5.00\cdot 10^{-3}$ $4.94\cdot 10^{-08}$ $7.1$ $1.39\cdot 10^{-07}$ $6.9$ $4.17\cdot 10^{-07}$ $6.9$
$50\cdot 10^{-3} 2.2$$21\cdot 10^{-10}$ $7.8$ $5.84\cdot 10^{-10}$ $7.9$ $2.63\cdot 10^{-09}$ $7.3$
$1.25\cdot 10^{-3}$ $1.31\cdot 10^{-12}$ $7.4$ $4.41\cdot 10^{-12}$ $7.0$ $2.06\cdot 10^{-11}$ $7.0$
Table 2: Convergence of the 2D-FC based solution of the nonlocal Poisson equation (9) in the settings of Example 3 with fixed $\delta=0.4$, fixed $M=d+1$, and two different values of $d$.
@@ 4.2 Accuracy and convergence test for nonlocal diffusion equation
A direct computation of the multipliers $m^{\delta,\beta}$ is provided in Table 3. We compute these multipliers using the implementation of the $_{2}F_{3}$ hypergeometric function provided by mpmath [johansson2013mpmath], as detailed in [alali2021fourier].
@@
$0.3$ 130.16228859689554-130.16228859689554 321.3202730766787-321.3202730766787 689.8419741563309-689.8419741563309
Table 3: The multipliers of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ applied on the function $u=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $r_{1}=r_{2}=15.6455$, with a fixed $\delta=0.3$ and three different values of $\beta$.
@@
We employ our proposed method to study the error convergence of the nonlocal diffusion problem (12) over the kite-shaped domain contained within the curve given by (14) using step sizes $h=0.02/2^{k},~k=0,1,2,3,4$ and time step size $\tau=5.0\cdot 10^{-07}$. We fix $\delta=0.3$, and choose three values of $\beta:$ $1.0$, $2.0$, and $2.5$, and we set the boundary function $b(x,y)=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)\exp({-2\pi^{2}\kappa t})$, where $\kappa=0.1$, $r_{1}=r_{2}=15.6455$, and the heat source, $s(x,y,t)$, is as given in (15).
@@
$00\cdot 10^{-3} 5.5$$53\cdot 10^{-09}$ $7.3$ $3.30\cdot 10^{-08}$ $7.5$ $3.23\cdot 10^{-07}$ $7.0$
@@
Table 4: Convergence of the 2D-FC based solution of the nonlocal diffusion equation (12) in the settings of Example 4 with fixed $\delta=0.3$, fixed $M=d+1$, two different values of $d$: 44 and $5$, and three choices of $\beta$: $1.0$, $2.0$, and $2.5$.
@@ 4.3 Study of discontinuities for the nonlocal Poisson equation
Let $\Omega$ be the kite-shaped domain contained within the curve defined in (14). We solve the nonlocal Poisson equation
$\begin{cases}\mathcal{L}^{\delta,\beta}u(x,y)=4,&(x,y)\in\Omega\\ u(x,y)=x^{2}+y^{2},&(x,y)\in\Omega_{I}\end{cases}$ (16)
with $\delta=0.4$, three different values of $\beta:$ $1.0$, $2.0$, $3.0$, and we choose the 2D-FC parameters $d=4$, $M=5$, and $h=k_{1}=k_{2}=0.0025$.
In this example, since for any $\delta>0$ and $\beta<4$, $\mathcal{L}^{\delta,\beta}(x^{2}+y^{2})=4$ for all $x\in\Omega$, then the unique solution for this problem is given explicitly by $u(x,y)=x^{2}+y^{2}$, which is obviously a continuous function. The numerical solution obtained using our solver is consistent with the above observation as illustrated in Figure 8.
@@
Let $\Omega$ be the unit disk. We solve the nonlocal Poisson equation
$\begin{cases}\mathcal{L}^{\delta,\beta}u(x,y)=|\sin(\pi x)\sin(\pi y)|,&(x,y)\in\Omega\\ u(x,y)=-x^{2},&(x,y)\in\Omega_{I}\end{cases}$ (17)
with $\delta=0.2$, four values of $\beta:$ $1.5$, $2.0$, $2.2$, $3.1$, and we choose the 2D-FC parameters $d=4$, $M=5$, and $h=k_{1}=k_{2}=0.0025$.
@@
$\text{magnitude of the jump}=\left|\widetilde{u}(x_{\theta},y_{\theta})-b(x_{\theta},y_{\theta})\right|,$
@@
Let $\Omega$ be the kite-shaped domain contained within the curve defined in (14) and let
@@
We solve the nonlocal Poisson equation
$\begin{cases}\mathcal{L}^{\delta,\beta}u(x,y)=f(x,y),&(x,y)\in\Omega\\ u(x,y)=x^{2}+y^{2},&(x,y)\in\Omega_{I}\end{cases}$ (18)
with $\delta=0.5$, four different values of $\beta:$ $1.2$, $2.0$, $2.5$, $3.1$, and we choose the 2D-FC parameters $d=4$, $M=5$, and $h=k_{1}=k_{2}=0.0025$.
@@
$\text{magnitude of the jump}=\left|\widetilde{u}_{\text{int}}(x_{\theta},y_{\theta})-\widetilde{u}_{\text{ext}}(x_{\theta},y_{\theta})\right|,$
@@
To make a connection between the numerical results in this work, in particular Example 7, and previous mathematical analysis in this area, we note that regularity results for the nonlocal Poisson equation on periodic domains have been developed in [alali2021fourier]. However, regularity of solutions for the nonlocal equation on bounded domains is still an open problem.
For the periodic case, Theorem 4.1 in [alali2021fourier] implies that if $f\in L^{2}$, then $u\in H^{\beta-n}$. In particular, in the case when when $n=2$ and $3<\beta<4$, then $u\in H^{\beta-2}$, which implies that $u$ is continuous using the embedding theorem. Using intuition from the periodic case combined with the numerical observations in Example 7, we expect that this smoothing behavior as $\beta$ increases would be true in the case of bounded domains as stated in the following conjecture.
@@
Let $u$ be the solution of (9) and assume that $f\in L^{2}(\Omega)$. Then, if $3<\beta<4$, then $u\in C(\Omega)$.
@@ 4.4 Evolution of discontinuities for nonlocal diffusion equation
where $r(\theta)=1.1+\cos(7\theta)/20+\sin(4\theta)/30,$ and $0\leq\theta\leq 2\pi$. We solve the nonlocal diffusion equation
$\begin{cases}u_{t}(x,y,t)=\mathcal{L}^{\delta,\beta}u(x,y,t),&(x,y,t)\in\Omega\times(0,T)\\ u(x,y,0)=\sin(2\pi x)\cos(2\pi y),&(x,y)\in\Omega\\ u(x,y,t)=-(1+x^{2}+y^{2}),&(x,y)\in\Omega_{I}\end{cases}$ (20)
with $\delta=0.2$ and two different values of $\beta$: $4.0$ and $1.5$ at four different time instances: $0.02$, $0.03$, $0.04$, and $0.05$ with time step size $\tau=2.5\cdot 10^{-7}$. We choose the 2D-FC parameters $d=4$, $M=5$, and $h=k_{1}=k_{2}=0.005$.
@@
Let $\Omega$ be the domain contained within the curve given in (19) and let
$u_{0}(x,y)=\begin{cases}2,&\text{ if }x^{2}+y^{2}<0.3,\\ \exp({-x^{4}-y^{4}}),&\text{ otherwise}.\end{cases}$
We solve the nonlocal diffusion equation
$\begin{cases}u_{t}(x,y,t)=\mathcal{L}^{\delta,\beta}u(x,y,t),&(x,y,t)\in\Omega\times(0,T)\\ u(x,y,0)=u_{0}(x,y),&(x,y)\in\Omega\\ u(x,y,t)=0,&(x,y)\in\Omega_{I}\end{cases}$ (22)
with $\delta=0.2$ and two values of $\beta$: $4.0$ and $1.0$ at four different time instances: $t=0.001$, $0.005$, $0.01$, and $0.03$ with time step size $\tau=2.5\cdot 10^{-7}$. We choose the 2D-FC parameters $d=4$, $M=5$, and $h=k_{1}=k_{2}=0.005$.
@@
5 Discussion
Fourier spectral methods are traditionally well-suited for periodic problems. In this work, we extend their applicability to nonlocal equations on bounded domains by constructing periodic extensions of the problem via Fourier continuation. This enables efficient computations while maintaining high-order convergence for smooth solutions and effectively addressing challenges associated with nonlocal operators near boundaries.
We also conduct numerical experiments to investigate solution regularity for two-dimensional nonlocal Poisson and diffusion equations on bounded domains, with particular attention to discontinuities. Our simulations reveal that for integrable kernels, discontinuities may develop in the solution. In contrast, for singular kernels with $\beta>n+1$, the numerical solution appears to approximate a continuous function within the interior of the domain–an observation formalized in Conjecture 1.
While our convergence analysis focuses on smooth solutions, we observe a decline in convergence order when discontinuities are present. Extending this approach to problems involving sharp interfaces or discontinuities, where limited regularity can reduce the effectiveness of spectral methods, remains an important direction for future research.